# Relational Graph Transformer

### v1

@@ 1 Introduction
Real-world enterprise data, such as financial transactions, supply chain data, e-commerce records, product catalogs, customer interactions, and electronic health records, are predominantly stored in relational databases [8]. These databases typically consist of multiple tables, each dedicated to different entity types, interconnected through primary and foreign key links. This abstraction underpins large quantities of complex, dynamically updated data that scale with business volume, storing potentially immense, unexploited knowledge [13]. However, extracting predictive patterns from such data has traditionally depended on manual feature engineering within complex machine learning pipelines, requiring the transformation of multi-table records into flat feature vectors suitable for models like deep neural networks and decision trees [6].
@@
Relational Deep Learning. To enable end-to-end deep learning, relational databases can be represented as relational entity graphs [13], where nodes correspond to entities and edges capture primary-foreign key relationships. This graph-based representation allows Graph Neural Networks (GNNs) to learn abstract features directly from the underlying data structure, effectively modeling complex dependencies for various downstream prediction tasks. With this setup, which is termed as Relational Deep Learning (RDL), GNNs reduce or eliminate the need for manual feature engineering and often lead to improved performance [42], at a fraction of the traditional model development cost.
@@
Existing gaps. Despite their effectiveness, standard message-passing GNN architectures [15, 29, 18, 48] have notable limitations, such as insufficient structural expressiveness [52, 37, 34] and restricted long-range modeling capabilities [1]. For example, consider an e-commerce database with three tables: customers, transactions, and products, which can be represented as a relational entity graph as in Figure 1. In a standard GNN, transactions are always two hops away from each other, connected only through shared customers. This creates an information bottleneck: transaction-to-transaction patterns require multiple layers of message passing, while product relationships remain entirely indirect in shallow networks. Furthermore, products would never directly interact in a two-layer GNN [42], as their messages must pass through both a transaction and a customer, highlighting the inherent structural constraints of GNN architectures that restrict capturing long-range dependencies.
@@
Graph Transformers (GTs) have emerged as more expressive models for graph learning, utilizing self-attention in the full graph to increase the range of information flow and additionally, incorporating positional and structural encodings (PEs/SEs) to better capture graph topology [9, 53, 41]. These advances have produced strong results across domains [38], including foundation models for molecular graphs [46]. However, many GT designs are limited to non-temporal, homogeneous, and small-scale graphs, assumptions that do not hold for relational entity graphs (REGs) [13], which are typically (i) heterogeneous, with different tables representing distinct node types; (ii) temporal, with entities often associated with timestamps and requiring careful handling to prevent data leakage; (iii) large-scale, containing millions or more records across multiple interconnected tables. In particular, existing PEs often require precomputation, depend on graph size, and typically do not scale well to large, heterogeneous, or dynamic graphs [3, 26]. For instance, node2vec [17], while more efficient than Laplacian or random walk PEs, can become prohibitively expensive and impractical to compute on massive graphs [40]. These limitations, along with the inability to capture the multi-dimensional complexity of relational structures, render current GTs inadequate for relational databases.
@@
  • Attention: We develop a transformer network that combines local and global representations, adapting existing GT architectures [41]. The model extracts features from the local tokens while simultaneously attending to learnable global tokens that act as soft centroids, effectively balancing fine-grained structural modeling with database-wide patterns [30].
  • Validation: We showcase RelGT’s effectiveness through a comprehensive evaluation on 21 tasks from RelBench [42]. RelGT consistently outperforms GNN baselines, with gains of up to 18%, establishing transformers as a powerful architecture for relational deep learning. Compared to HGT [20], a strong GT baseline for heterogeneous graphs, RelGT achieves better results without added computational cost, even when HGT uses Laplacian eigenvectors for positional encoding.
@@ 2.1 Relational Deep Learning
Relational Deep Learning (RDL) is an end-to-end representation learning framework that converts relational databases into graph structures, enabling neural networks to be applied directly and eliminating the need for manual feature extraction in multi-table data pipelines [13].
Definitions. Formally, we can define a relational database as the tuple $(T,R)$ comprising a collection of tables $T=\{T_{1},\ldots,T_{n}\}$ connected through inter-table relationships $R\subseteq T\times T$. A link $(T_{\text{fkey}},T_{\text{pkey}})\in R$ denotes a foreign key in one table referencing a primary key in another. Each table contains entities (rows) $\{v_{1},\ldots,v_{n_{T}}\}$, with each entity typically consisting of: (1) a unique identifier (primary key), (2) references to other entities (foreign keys), (3) entity-specific attributes, and (4) timestamp information indicating when the entity was created or modified. The structure of relational databases inherently forms a graph representation, called as relational entity graphs (REGs). An REG is formally defined as a heterogeneous temporal graph $G=(V,E,\phi,\psi,\tau)$, where nodes $V$ represent entities from the database tables, edges $E$ represent primary-foreign key relationships, $\phi$ maps nodes to their respective types based on source tables, $\psi$ assigns relation types to edges, and $\tau$ captures the temporal dimension through timestamps [13].
Challenges. Relational entity graphs exhibit three distinctive properties that set them apart from conventional graph data. First, their structure is fundamentally schema-defined, with topology shaped by primary-foreign key relationships rather than arbitrary connections, creating specific patterns of information flow that require specialized modeling approaches. Second, they incorporate temporal dynamics, as relational databases track events and interactions over time, necessitating techniques like time-aware neighbor sampling to prevent future information from leaking into past predictions. Third, they display multi-type heterogeneity, as different tables correspond to different entity types with diverse attribute schemas and data modalities, presenting challenges in creating unified representations that effectively integrate information across diverse node and edge types [44, 49]. These characteristics create both challenges and opportunities for GNN architectures, requiring models that can simultaneously address temporal evolution, heterogeneous information, and schema-constrained structures while processing potentially massive multi-table datasets.
@@ 2.2 RDL Methods
The baseline GNN approach introduced by [42] for RDL uses a heterogeneous GraphSAGE [18] model with temporal-aware neighbor sampling, which demonstrates significant improvements compared to traditional tabular methods like LightGBM [27] across all tasks in the RelBench benchmark. This baseline architecture leverages PyTorch Frame’s multi-modal feature encoders [19] to transform diverse entity attributes into initial feature embeddings that serve as input to the GNN. Several specialized architectures have been developed to address specific challenges in relational entity graphs. RelGNN [5] introduces composite message-passing with atomic routes to facilitate direct information exchange between neighbors of bridge and hub nodes, commonly found in relational structures. Similarly, ContextGNN [55] employs a hybrid approach, combining pair-wise and two-tower representations, specifically optimized for recommendation tasks in RelBench.
Beyond pure GNN approaches, retrieval-augmented generation techniques [51] and hybrid tabular-GNN methods [32] have also demonstrated comparable or slightly superior performance to the standard GNN baseline, while showing the use of LLMs [16] and inference speedups, respectively. These approaches confirm the effectiveness of graph, tabular, and LLM-based methods for downstream predictions in relational databases. However, these methods typically optimize specific aspects of the problem, failing to incorporate broader advances from GTs in general graph learning.
@@ 2.3 Graph Transformers
Graph Transformers extend the self-attention mechanism from sequence modeling [47] to graph-structured data, offering powerful alternatives to traditional GNNs [9]. These models typically restrict attention to local neighborhoods, functioning as message-passing networks with attention-based aggregation [24, 2], while positional encodings are developed based on Laplacian eigenvectors [10]. Subsequent Graph Transformers incorporate global attention mechanisms, allowing all nodes to attend to one another [53, 36, 31]. This moves beyond the local neighborhood limitations of standard GNNs [1], albeit at the cost of significantly increased computational complexity.
Modern GT architectures have improved the aforementioned early works by creating effective structural encodings and ensuring scalability to medium and large-scale graphs. For structural expressiveness of the node tokens, several positional and structural encoding methods have been developed [12, 3, 33, 22, 26] to inject the input graph topology. For scalability, various strategies have emerged including hierarchical clustering that coarsens graphs [57, 59], sparse attention mechanisms that reduce computational cost [41, 45], and neighborhood sampling techniques for processing massive graphs [58, 4, 30, 11]. Models like GraphGPS [41] combine these advances through hybrid local-global designs that maintain Transformers’ global context advantages while ensuring practical efficiency when scaling to medium and large graph datasets. However, these approaches exhibit several key limitations: they are largely confined to static graphs, and lack mechanisms to handle multiple node and edge types. While specialized Transformers for heterogeneous graphs exist [20, 35, 59, 56], integrating them, alongside other aforementioned methods, into the RDL pipeline remains challenging. This is primarily because adapting positional encodings under precomputation constraints is difficult, compounded by the complexity of modeling large-scale, temporal, and heterogeneous relational entity graphs (REGs).
@@ 3.1 Tokenization
Traditional Transformers in natural language processing represent text through tokens with two primary elements: (i) token identifiers (or features) that denotes the token from a vocabulary set and (ii) positional encodings that represent sequential structure [47]. For example, a token can correspond to a word and its positional encoding can correspond to its order in the input sentence. Similarly, Graph Transformers generally adapt this two-element representation to graphs, where nodes are tokens with features, and graph positional encodings provide structural information [9, 28, 41]. Although this two-element approach works well for homogeneous static graphs, it becomes computationally inefficient when trying to encode multiple aspects of graph structural information for REGs.
@@ 3.1.1 Proposed Approach
[…]
@@
1. Node Feature Encoder. The node features $x_{v_{j}}$, representing the columnar attributes of the node $v_{j}$ in REG (which corresponds to a table row in a database), are encoded into a $d$-dimensional embedding. Each modality, such as numerical, categorical, multi-categorical, text, and image data, is encoded separately using modality-specific encoders following [19], and the resulting representations are then aggregated into a unified $d$-dimensional embedding.
@@
[…]
@@
5. Subgraph PE Encoder. Finally, for capturing local graph structure that can otherwise not be represented by other elements of the token, we apply a light-weight GNN to the subgraph. This GNN encoder effectively preserves important structural relationships, such as complex cycles and quasi-cliques between entities [25], as well as parent-child relationships (e.g., a product node within the local subgraph corresponding to specific transactions), and can be written as:
@@
One key advantage of using random node features in this GNN encoder is that it breaks structural symmetries between the subgraph topology and node attributes, thereby increasing the expressive power of GNN layers [43]. However, a fixed random initialization would destroy permutation equivariance which is a critical property for generalization. To address this challenge, we resample $Z_{\text{random}}$ independently at every training step. This ‘stochastic initialization’ approach can be viewed as a relaxed version of the learnable PE method described in [26], thus approximately preserving permutation equivariance while retaining the expressivity gains afforded by the randomization.
@@
${h}_{\text{token}}(v_{j})=O\cdot\left[{h}_{\text{feat}}(v_{j})\,||\,{h}_{\text{type}}(v_{j})\,||\,{h}_{\text{hop}}(v_{i},v_{j})\,||\,{h}_{\text{time}}(v_{i},v_{j})\,||\,{h}_{\text{pe}}(v_{j})\right]$ (6)
where ||||| | denotes the concatenation of the individual encoder outputs, and $O\in\mathbb{R}^{5d\times d}$ is a learnable matrix to mix the embeddings. This multi-element approach provides a comprehensive token representation that explicitly captures node features, type information, structural position, temporal dynamics, and local topology without requiring expensive computation on the graph structure.

### v2

@@ 1 Introduction
Real-world enterprise data, such as financial transactions, supply chain data, e-commerce records, product catalogs, customer interactions, and electronic health records, are predominantly stored in relational databases (Codd, 1970). These databases typically consist of multiple tables, each dedicated to different entity types, interconnected through primary-foreign key links. This abstraction underpins large quantities of complex, dynamically updated data that scale with business volume, storing potentially immense, unexploited knowledge (Fey et al., 2024). However, extracting predictive patterns from such data has traditionally depended on manual feature engineering within complex machine learning pipelines, requiring the transformation of multi-table records into flat feature vectors suitable for models like deep neural networks and decision trees (Chen & Guestrin, 2016).
@@
Relational Deep Learning. To enable end-to-end deep learning, relational databases can be represented as relational entity graphs (Fey et al., 2024), where nodes correspond to entities and edges capture primary-foreign key relationships. This graph-based representation allows Graph Neural Networks (GNNs) to learn abstract features directly from the underlying data structure, effectively modeling complex dependencies for various downstream prediction tasks. With this setup, which is termed as Relational Deep Learning (RDL), GNNs reduce or eliminate the need for manual feature engineering and often lead to better performance (Robinson et al., 2024), at a fraction of the traditional model development cost.
@@
Existing gaps. Despite their effectiveness, standard message-passing GNN architectures (Gilmer et al., 2017; Kipf & Welling, 2016; Hamilton et al., 2017; Velickovic et al., 2017) have notable limitations, such as insufficient structural expressiveness (Xu et al., 2019; Morris et al., 2019; Loukas, 2019) and restricted long-range modeling capabilities (Alon & Yahav, 2020). For example, consider an e-commerce database with three tables: customers, transactions, and products, which can be represented as a relational entity graph as in Figure 1. In a standard GNN, transactions are always two hops away from each other, connected only through shared customers. This creates an information bottleneck: transaction-to-transaction patterns require multiple layers of message passing, while product relationships remain entirely indirect in shallow networks. Furthermore, products would never directly interact in a two-layer GNN (Robinson et al., 2024), as their messages must pass through both a transaction and a customer, highlighting the inherent structural constraints of GNN architectures that restrict capturing long-range dependencies.
@@
Graph Transformers (GTs) have emerged as more expressive models for graph learning, utilizing self-attention in the full graph to increase the range of information flow and additionally, incorporating positional and structural encodings (PEs/SEs) to better capture graph topology (Dwivedi & Bresson, 2021; Ying et al., 2021; Rampášek et al., 2022). These advances have produced strong results across domains (Müller et al., 2023), including foundation models for molecular graphs (Sypetkowski et al., 2024). However, many GT designs are limited to non-temporal, homogeneous, and small-scale graphs, assumptions that do not hold for relational entity graphs (REGs) (Fey et al., 2024), which are typically (i) heterogeneous, with different tables representing distinct node types; (ii) temporal, with entities often associated with timestamps and requiring careful handling to prevent data leakage; (iii) large-scale, containing millions or more records across multiple interconnected tables. In particular, existing PEs often require precomputation, depend on graph size, and typically do not scale well to large, heterogeneous, or dynamic graphs (Cantürk et al., 2023; Kanatsoulis et al., 2025). For instance, node2vec (Grover & Leskovec, 2016), while more efficient than Laplacian or random walk PEs, can become prohibitively expensive and impractical to compute on massive graphs (Postăvaru et al., 2020). These limitations, along with the inability to capture the multi-dimensional complexity of relational structures, render current GTs inadequate for relational databases.
@@
  • Attention: We develop a transformer network that combines local and global representations, adapting existing GT architectures (Rampášek et al., 2022). The model extracts features from the local tokens while simultaneously attending to learnable global tokens that act as soft centroids, effectively balancing fine-grained structural modeling with database-wide patterns.
  • Validation: We showcase RelGT’s effectiveness through a comprehensive evaluation on 21 tasks from RelBench (Robinson et al., 2024). RelGT consistently outperforms GNN baselines, with gains of up to 18%, establishing transformers as a powerful architecture for relational deep learning. Compared to HGT, a strong GT baseline for heterogeneous graphs, RelGT achieves better results without added computational cost, even when HGT uses Laplacian eigenvectors for PE.
@@ 2.1 Relational Deep Learning
Relational Deep Learning is an end-to-end learning framework that converts relational databases into graph structures, enabling direct use of GNNs for representation learning (Fey et al., 2024).
Definitions. Formally, we can define a relational database as the tuple $(T,R)$ comprising a collection of tables $T=\{T_{1},\ldots,T_{n}\}$ connected through inter-table relationships $R\subseteq T\times T$. A link $(T_{\text{fkey}},T_{\text{pkey}})\in R$ denotes a foreign key in one table referencing a primary key in another. Each table contains entities (rows) $\{v_{1},\ldots,v_{n_{T}}\}$, with each entity typically consisting of: (1) a unique identifier (primary key), (2) references to other entities (foreign keys), (3) entity-specific attributes, and (4) timestamp information indicating when the entity was created or modified. The structure of relational databases inherently forms a graph representation, called as relational entity graphs (REGs). An REG is formally defined as a heterogeneous temporal graph $G=(V,E,\phi,\psi,\tau)$, where nodes $V$ represent entities from the database tables, edges $E$ represent primary-foreign key relationships, $\phi$ maps nodes to their respective types based on source tables, $\psi$ assigns relation types to edges, and $\tau$ captures the temporal dimension through timestamps (Fey et al., 2024).
Challenges. Relational entity graphs exhibit three distinctive properties that set them apart from conventional graph data. First, their structure is fundamentally schema-defined, with topology shaped by primary-foreign keys rather than arbitrary connections, creating specific patterns of information flow that require specialized modeling approaches. Second, they incorporate temporal dynamics, as relational databases track events and interactions over time, necessitating techniques like time-aware neighbor sampling to prevent future information from leaking into past predictions. Third, they display multi-type heterogeneity, as different tables correspond to different entity types with diverse attribute schemas and data modalities, presenting challenges in creating unified representations that effectively integrate information across diverse node and edge types (Schlichtkrull et al., 2018; Wang et al., 2019). These characteristics create both challenges and opportunities for GNN architectures, requiring models that can simultaneously address temporal evolution, heterogeneous information, and schema-constrained structures while processing potentially massive multi-table datasets.
@@ 2.2 RDL Methods
The baseline GNN approach introduced by (Robinson et al., 2024) for RDL uses a heterogeneous GraphSAGE (Hamilton et al., 2017) model with temporal-aware neighbor sampling, which demonstrates significant improvements compared to traditional tabular methods like LightGBM (Ke et al., 2017) across most of the tasks in the RelBench benchmark. This baseline architecture leverages PyTorch Frame’s multi-modal feature encoders (Hu et al., 2024) to transform diverse entity attributes into initial feature embeddings that serve as input to the GNN. Several specialized architectures have been developed to address specific challenges in relational entity graphs. RelGNN (Chen et al., 2025) introduces composite message-passing with atomic routes to facilitate direct information exchange between neighbors of bridge and hub nodes, commonly found in relational structures. Similarly, ContextGNN (Yuan et al., 2024) employs a hybrid approach, combining pair-wise and two-tower representations, specifically optimized for recommendation tasks in RelBench.
Beyond pure GNN approaches, retrieval-augmented generation techniques (Wydmuch et al., 2024) and hybrid tabular-GNN methods (Lachi et al., 2024) have also demonstrated comparable or superior performance to the standard GNN baseline, while showing the use of LLMs (Grattafiori et al., 2024) and inference speedups, respectively. These approaches confirm the effectiveness of graph, tabular, and LLM-based methods for downstream predictions in RDL. However, these methods typically optimize specific aspects of the problem, failing to incorporate broader advances from GTs in general.
@@ 2.3 Graph Transformers
Graph Transformers extend the self-attention mechanism from sequence modeling (Vaswani et al., 2017) to graph-structured data, offering powerful alternatives to traditional GNNs (Dwivedi & Bresson, 2021). These models typically restrict attention to local neighborhoods, functioning as message-passing networks with attention-based aggregation (Joshi, 2020; Bronstein et al., 2021), while positional encodings are developed based on Laplacian eigenvectors (Dwivedi et al., 2020). Subsequent Graph Transformers incorporate global attention mechanisms, allowing all nodes to attend to one another (Ying et al., 2021; Mialon et al., 2021; Kreuzer et al., 2021). This moves beyond the local neighborhood limitations of standard GNNs (Alon & Yahav, 2020), albeit at the cost of significantly increased computational complexity.
Modern GT architectures have improved the aforementioned early works by creating effective structural encodings and ensuring scalability to medium and large-scale graphs. For structural expressiveness of the node tokens, several positional and structural encoding methods have been developed (Dwivedi et al., 2022; Cantürk et al., 2023; Lim et al., 2022; Huang et al., ; Kanatsoulis et al., 2025) to inject the input graph topology. For scalability, various strategies have emerged including hierarchical clustering that coarsens graphs (Zhang et al., 2022; Zhu et al., 2023), sparse attention mechanisms that reduce computational cost (Rampášek et al., 2022; Shirzad et al., 2023), and neighborhood sampling techniques for processing massive graphs (Zhao et al., 2021; Chen et al., 2022; Dwivedi et al., 2023). Models like GraphGPS (Rampášek et al., 2022) combine these advances through hybrid local-global designs that maintain Transformers’ global context advantages while ensuring practical efficiency when scaling to medium and large graph datasets. However, these approaches exhibit several key limitations: they are largely confined to static graphs, and lack mechanisms to handle multiple node and edge types. While specialized Transformers for heterogeneous graphs exist (Hu et al., 2020; Mao et al., 2023; Zhu et al., 2023; Zhai et al., 2024), integrating them, alongside other aforementioned methods, into the RDL pipeline remains challenging. This is primarily because adapting PEs under precomputation constraints is difficult, compounded by the complexity of modeling large-scale, temporal, and heterogeneous relational entity graphs (REGs).
2.4 Addressing Challenges
While heterogeneous graph transformers (Hu et al., 2020) and temporal graph methods exist, no prior GT architecture effectively handles relational entity graphs where heterogeneity, temporality, and rich entity attributes co-occur within schema-defined database structures. Heterogeneous Temporal Graph Transformers like HTGformer (Wang, 2025) process heterogeneity and temporality through separate, iterative modules without graph positional encodings, a component now considered essential in modern GTs (Rampášek et al., 2022; Müller et al., 2023), and do not address the multimodal attributes or schema constraints inherent to relational databases (Hu et al., 2024). Existing subgraph-based GTs (Zhao et al., 2021; Chen et al., 2022) focus on scalability for homogeneous graphs without mechanisms for heterogeneity or temporal dynamics.
We address these limitations by recognizing that relational entity graphs require a rethinking on how their multimodal attributes and comprehensive graph structure are jointly processed. We systematically decompose the information coming from the REGs into specialized components that can be independently learned and composed, as we describe in detail in Section 3. This principled design enables GTs to handle the unique combination of heterogeneity, temporality, and schema-defined structures in relational databases without expensive global precomputation.
@@ 3.1 Tokenization
Traditional Transformers in NLP represent text through tokens with two primary elements: (i) token identifiers (or features) that denotes the token from a vocabulary set and (ii) positional encodings that represent sequential structure (Vaswani et al., 2017). For example, a token can correspond to a word and its positional encoding can correspond to its order in the input sentence. Similarly, Graph Transformers generally adapt this two-element representation to graphs, where nodes are tokens with features, and graph positional encodings provide structural information. Although this two-element approach works well for homogeneous static graphs, it becomes computationally inefficient when trying to encode multiple aspects of graph structural information for REGs.
@@ 3.1.1 Proposed Approach
Sampling and token elements. The tokenization process in RelGT converts a REG $G=(V,E,\phi,\psi,\tau)$ into sets of tokens suitable for processing by the Transformer network. Specifically, as shown in Fig 2, for each training seed node $v_{i}\in V$, we first sample a fixed set of $K$ neighboring nodes $v_{j}$ from within 2 hops of the local neighborhood using temporal-aware sampling [^2] When fewer than $K$ neighbors are available within 2 hops, we use randomly selected nodes as fallback tokens to maintain the fixed size $K$, ensuring consistent computational complexity regardless of local structure., ensuring that only nodes with timestamps $\tau(v_{j})\leq\tau(v_{i})$ are included to prevent temporal leakage. Each token in this set is represented by a 5-tuple: $(x_{v_{j}},\phi(v_{j}),p(v_{i},v_{j}),\tau(v_{j})-\tau(v_{i}),\text{GNN-PE}_{v_{j}})$, where, (i) node feature ($x_{v_{j}}$) denotes the raw features derived from entity attributes in the database, (ii) node type ($\phi(v_{j})$) is a categorical identifier corresponding to the entity’s originating table, (iii) relative hop distance ($p(v_{i},v_{j})$) captures the structural distance between the seed node $v_{i}$ and the neighbor node $v_{j}$, (iv) relative time ($\tau(v_{j})-\tau(v_{i})$) represents the temporal difference between the neighbor and seed node, and (v) finally, subgraph based PE ($\text{GNN-PE}_{v_{j}}$) provides a graph PE for each node within the sampled subgraph, generated by applying a lightweight GNN to the subgraph’s adjacency matrix with random node feature initialization (Sato et al., 2021; Kanatsoulis et al., 2025).
@@
1. Node Feature Encoder. The node features $x_{v_{j}}$, representing the columnar attributes of the node $v_{j}$ in REG (which corresponds to a table row in a database), are encoded into a $d$-dimensional embedding. Each modality, such as numerical, categorical, multi-categorical, text, and image data, is encoded separately using modality-specific encoders following (Hu et al., 2024), and the resulting representations are then aggregated into a unified $d$-dimensional embedding.
@@
where $\text{MultiModalEncoder}(\cdot)$ is unified feature encoder adapted from (Hu et al., 2024).
@@
5. Subgraph PE Encoder. Finally, for capturing local graph structure that can otherwise not be represented by other token elements, we apply a light-weight GNN to the subgraph. This GNN encoder effectively preserves important structural relationships, such as complex cycles and quasi-cliques between entities (Kanatsoulis & Ribeiro, 2024), as well as parent-child relationships (e.g., a product node within the local subgraph corresponding to specific transactions), and can be written as:
@@
One key advantage of using random node features in this GNN encoder is that it breaks structural symmetries between the subgraph topology and node attributes, thereby increasing the expressiveness of GNN layers (Sato et al., 2021). However, a fixed random initialization would destroy permutation equivariance, a critical property for generalization. To address this, we resample $Z_{\text{random}}$ independently at every training step. This ‘stochastic initialization’ approach can be viewed as a relaxed version of the learnable PE method described in Kanatsoulis et al. (2025), thus approximately preserving permutation equivariance while retaining the expressivity gains afforded by the randomization.
@@
[…]
where $||$ denotes the concatenation of the individual encoder outputs, and $O\in\mathbb{R}^{5d\times d}$ is a learnable matrix to mix the embeddings. This multi-element approach provides a comprehensive token representation that explicitly captures node features, type information, structural position, temporal dynamics, and local topology without requiring expensive computation on the graph structure.